OpenStudy (anonymous):

Amanda's poster is 0.85 meter long. Orlando's poster is 0.55 meter less than 1 meter long. Which expression shows the difference in the lengths of the posters?

OpenStudy (jamanorthi):

is there an option?

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) are there options for this? Orlando's poster would equal (1- 0.55) meters Amanda's poster would equal 0.85 meters the difference between the two lengths would be represented as 0.85 - (1-0.55) = ?

OpenStudy (anonymous):

1-0.55=orlando's poster. then, difference between Orlando's poster and Amanda is o.85-(1-0.55)=0.85-0.45

OpenStudy (anonymous):

the answer would be 0.4 because u subtract 1 from 0.55 and u will 0.45 and then u will take 0.45 from 0.85 and will 0.4!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yea, thats the same as i said above ... lol,"= 0.85-0.45"